OpenStudy (anonymous):

Factor by grouping. y^2-4y-6y+24

OpenStudy (xapproachesinfinity):

factor by grouping \[y^2-4y-6y+24=(y^2-4y)-(6y-24)=y(y-4)-6(y-4)\]

OpenStudy (xapproachesinfinity):

can you continue

OpenStudy (anonymous):

I don't know how

OpenStudy (xapproachesinfinity):

is there a common factor?

OpenStudy (xapproachesinfinity):

is not (y-4) common?

OpenStudy (xapproachesinfinity):

isn't it a common factor?

OpenStudy (anonymous):

Yes it is

OpenStudy (xapproachesinfinity):

do you know what a common factor mean 45-9 what is the common factor?

OpenStudy (anonymous):

It means the same numbers. I know what it means. I am on my phone and can't see the whole responses.

OpenStudy (xapproachesinfinity):

btw 45 nd 9 what is the common number?